# Communication Optimization for Decentralized Learning atop Bandwidth-limited Edge Networks

### v1

@@ I Introduction
As explained later in Section II-C, only agent pairs corresponding non-zero entries in the mixing matrix need to communicate during DFL, allowing the mixing matrix to control the communication demands between agents. This observation has triggered a series of studies on how to optimally design the mixing matrix such as [8, 9, 6, 10, 11]. However, most existing works made the simplistic assumption that each pair of logically adjacent agents are also physically adjacent in the underlying communication network. This assumption has led to simplistic models of the communication cost under a given design, e.g., the maximum degree [10, 11], the minimum number of matchings [8, 9], or other functions of the communication graph [6]. The underlying assumption of all these cost models is that the communication cost (e.g., time) for a given agent-to-agent communication graph only depends on the topology of this graph. However, this assumption may not hold in practice, as agents typically communicate through an underlying communication network and the cost for a given set of communications also depends on the state (e.g., topology, routing, link capacities) of this network. In this work, we study the communication optimization for such overlay-based DFL, with focus on DFL over bandwidth-limited edge networks.
@@ I-A Related Work
Communication cost reduction. There are two general approaches for reducing the communication cost in FL: reducing the amount of data per communication through compression, e.g., [5], and reducing the number of communications, e.g., [20, 21, 22]. The two approaches can be combined for further improvement [23, 7]. Instead of either activating all the links or activating none, it has been recognized that better efficiency can be achieved by activating subsets of links. To this end, [23, 7] proposed an event-triggered mechanism and [8, 9, 6] proposed to activate links with predetermined probabilities. In this regard, our work designs predetermined link activation as in [8, 9, 6], which provides more predictable performance than event-triggered mechanisms, but we consider a cost model tailored to overlay-based DFL: instead of measuring the communication time by the number of matchings [8, 9, 6] or the maximum degree [10, 11], we use the minimum time to complete the activated agent-to-agent communications over a bandwidth-limited underlay, while taking into account limited link capacities and possibly shared links.
Topology design in DFL. The logical topology defining the neighborhoods of learning agents is an important design parameter in DFL. The impact of this topology on the convergence rate of DFL has been mostly captured through the spectral gap of the mixing matrix [1, 24, 25, 26, 27] or equivalent parameters [8, 9]. Although recent works have identified other parameters that can impact the convergence rate, such as the effective number of neighbors [28] and the neighborhood heterogeneity [11], these results just pointed out additional factors and did not invalidate the impact of spectral gap. Based on the identified convergence parameters, several solutions have been proposed to design the logical topology to balance the convergence rate and the cost per communication round [8, 9, 6, 11], and some solutions combined topology design with other optimizations (e.g., bandwidth allocation [29], model pruning [27]) for further improvement. Our work also includes topology design based on a parameter related to the spectral gap, but we consider the joint optimization of the topology and the routing within the overlay in an overlay-underlay network.
@@ III-B1 Link Weight Design
[32, Theorem 2] Under assumptions (1)–(3), if the mixing matrix $\bm{W}$, which is symmetric with each row/column summing to one, satisfies that $\rho:=\|\bm{W}-\bm{J}\|<1$, then D-PSGD can achieve $\frac{1}{K}\sum_{k=1}^{K}\mathbbm{E}[\|\nabla F(\bm{\overline{x}}^{k})\|^{2}]\leq\epsilon$ for any given $\epsilon>0$ ($\overline{\bm{x}}^{(k)}:={1\over m}\sum_{i=1}^{m}\bm{x}^{(k)}_{i}$) when the number of iterations reaches
@@
Remark: The original version of [32, Theorem 2] covers more general cases where the mixing matrices can be random and time-varying. However, for the tractability of design, we will focus on the case of a single deterministic mixing matrix, in which case the bound in [32, Theorem 2] is reduced to (13) as shown in [18]. While there exist other convergence bounds for D-PSGD such as [8, 28, 26, 11], we choose Theorem III.3 as the theoretical foundation of our design due to the generality of its assumptions as explained in [32].
@@ IV-A1 Dataset and ML Model
We train a ResNet-50 model with 23,616,394 parameters and a model size of 94.47 MB (under precision FP32) for image classification on the CIFAR-10 dataset, which contains 60,000 color images divided into 10 classes. We use 50,000 images for training and the remaining 10,000 images for testing. The dataset undergoes standard preprocessing, including normalization and one-hot encoding of the labels. We use a mini‐batch size of 64 and an adaptive learning rate that is 0.1 for the first 30 epochs, 0.05 for the next 30 epochs, and 0.01 thereafter. These settings are sufficient for D-PSGD to achieve convergence under all the evaluated designs.

### v2

@@ I Introduction
As explained later in Section II-C, only agent pairs corresponding non-zero entries in the mixing matrix need to communicate during DFL, allowing the mixing matrix to control the communication demands between agents. This observation has triggered a series of studies on how to optimally design the mixing matrix such as [8, 6, 9, 10]. However, most existing works made the simplistic assumption that each pair of logically adjacent agents are also physically adjacent in the underlying communication network. This assumption has led to simplistic models of the communication cost under a given design, e.g., the maximum degree [9, 10], the minimum number of matchings [8], or other functions of the communication graph [6]. The underlying assumption of all these cost models is that the communication cost (e.g., time) for a given agent-to-agent communication graph only depends on the topology of this graph. However, this assumption may not hold in practice, as agents typically communicate through an underlying communication network and the cost for a given set of communications also depends on the state (e.g., topology, routing, link capacities) of this network. In this work, we study the communication optimization for such overlay-based DFL over bandwidth-limited edge networks.
@@ I-A Related Work
Communication cost reduction. There are two general approaches for reducing the communication cost in FL: reducing the amount of data per communication through compression, e.g., [5], and reducing the number of communications, e.g., [19, 20]. The two approaches can be combined for further improvement [21, 7]. Instead of either activating all the links or activating none, it has been recognized that better efficiency can be achieved by activating subsets of links. To this end, [21, 7] proposed an event-triggered mechanism and [8, 6] proposed to activate links with predetermined probabilities. In this regard, our work designs predetermined link activation as in [8, 6], which provides more predictable performance than event-triggered mechanisms, but we consider a cost model tailored to overlay-based DFL: instead of measuring the communication time by the number of matchings [8, 6] or the maximum degree [9, 10], we use the actual time to complete the activated agent-to-agent communications over a bandwidth-limited underlay with possibly shared links.
Topology design in DFL. The logical topology defining the neighborhoods of learning agents is an important design parameter in DFL. The impact of this topology on the convergence rate of DFL has been mostly captured through the spectral gap of the mixing matrix [1, 22, 23, 24, 25] or equivalent parameters [8]. Although recent works have identified other parameters that can impact the convergence rate, such as the effective number of neighbors [26] and the neighborhood heterogeneity [10], these results just pointed out additional factors and did not invalidate the impact of spectral gap. Based on the identified convergence parameters, several solutions have been proposed to design the logical topology to balance the convergence rate and the cost per communication round [8, 6, 10], and some solutions combined topology design with other optimizations (e.g., bandwidth allocation [27], model pruning [25]) for further improvement. Our work also includes topology design based on a parameter related to the spectral gap, but we consider the joint optimization of the topology and the routing within the overlay in an overlay-underlay network.
@@ III-B1 Link Weight Design
[30, Theorem 2] Under assumptions (1)–(3), if the mixing matrix $\bm{W}$, which is symmetric with each row/column summing to one, satisfies that $\rho:=\|\bm{W}-\bm{J}\|<1$, then D-PSGD can achieve $\frac{1}{K}\sum_{k=1}^{K}\mathbbm{E}[\|\nabla F(\bm{\overline{x}}^{k})\|^{2}]\leq\epsilon$ for any given $\epsilon>0$ ($\overline{\bm{x}}^{(k)}:={1\over m}\sum_{i=1}^{m}\bm{x}^{(k)}_{i}$) when the number of iterations reaches
@@
Remark: The original version of [30, Theorem 2] covers more general cases where the mixing matrices can be random and time-varying. However, for the tractability of design, we will focus on the case of a single deterministic mixing matrix, in which case the bound in [30, Theorem 2] is reduced to (13) as shown in [17]. While there exist other convergence bounds for D-PSGD such as [8, 26, 24, 10], we choose Theorem III.3 as the theoretical foundation of our design due to the generality of its assumptions as explained in [30].
@@ IV-A1 Dataset and ML Model
We train a ResNet-50 model with 23,616,394 parameters and a model size of 94.47 MB (under precision FP32) for image classification on the CIFAR-10 dataset, which contains 60,000 color images divided into 10 classes. We use 50,000 images for training and the remaining 10,000 images for testing. The dataset undergoes standard preprocessing, including normalization and one-hot encoding of the labels. We use a mini‐batch size of 64 and an adaptive learning rate that is 0.1 for the first 30 epochs, 0.05 for the next 30 epochs, and 0.01 thereafter. These settings are sufficient for D-PSGD to achieve convergence under all the evaluated designs. To check the generalizability of our results, we also train a 4-layer CNN model with 0.2 learning rate, which has 582,026 parameters and a model size of 2.33 MB, for digit recognition on the MNIST dataset, which contains 60,000 training images and 10,000 testing images. This model setup is based on the approach from [2]. We defer the results on MNIST to Appendix -B as the observations are similar.